# Incorporating Procedural Fairness in Flag Submissions on Social Media Platforms

### v1

@@ 4.1.2.   Flagging Scenarios
In summary, Figure 4 illustrates the flagging components and an example of flagging design flow a participant could have encountered in the survey. The left column enumerates each component comprising the scenario, specifically the independent variables and their levels. The right column shows one example of the flagging mechanism scenario a participant might have encountered. Once a participant selected one of the seven examples of inappropriate posts listed above, the flagging process (scenario) started. In this particular scenario, a participant would have been required to choose one of the rule violations from the primary classification scheme for the post they selected. Then, the participant would have been offered a text box to describe the rule violation in detail. The final step would have been a message of gratitude for flagging, informing them that an auto-moderator will review the post they flagged.
@@ 4.1.3.   Dependent Variables: Procedural Fairness
We used responses to these questions to operationalize participants’ perceptions of consistency, transparency, and voice, respectively. Participants indicated their responses to these questions on a 7-point Likert scale that ranged from Strongly Disagree to Strongly Agree. Additionally, the survey featured three open-ended questions asking participants to provide suggestions on how the flagging process could be further enhanced. These questions were:
@@ 4.1.4.   Usability
    “How demanding was it for you to use this flagging mechanism to report the post you selected?” (rated on the scale of ‘Very undemanding’ to ‘Very demanding’)
@@
    “How likely are you to use this flagging mechanism to report any inappropriate post on a social media platform you use?” (rated on the scale of ‘Very unlikely’ to ‘Very likely’.)
@@ 4.2. Deployment of the Survey
Our study was considered exempt from review by our institute’s [^5] We will reveal the institute name after the peer review is completed. IRB on November 16th, 2023. Before launching the survey, we conducted a pilot survey from May 1 to May 8 to gather feedback on the survey’s wording and organization. Based on this feedback, we revised the survey questionnaire by adjusting certain words and adding signposts. After this, we rolled out our main survey, which targeted U.S. adult internet users aged 18 and above. Participants were recruited through Lucid Theorem (https://lucidtheorem.com), a survey company providing access to nationally representative samples, from May 16 to May 18, 2024. Participants who completed the survey received \$1.50 in compensation via the Lucid system. In total, we collected responses from 3,154 participants.
We excluded respondents who opted out during the survey (N = 496), spent less than 1 minute or more than 50 minutes on the survey (N = 40), and exhibited straight-lining behavior by selecting the same response for all questions (N = 178). Following these pre-processing steps, we retained 2,936 survey responses, which we used for our subsequent analyses. Appendix A describes the demographic distribution of our survey sample. As mentioned above, the survey began with participants selecting the most inappropriate post (in their judgment) from the list of rule-violating posts presented in Table 2. Descriptive analysis showed the frequency distribution of choices made by survey participants to be as follows: Post 1 (21.2%), Post 2 (9.3%), Post 3 (23.6%), Post 4 (5.6%), Post 5 (6.1%), Post 6 (4.6%), and Post 7 (29.7%).
@@ 4.3.1.   Quantitative Analysis
We used SPSS Version 29 to analyze our quantitative data. Sec. 5.1 presents the results of our quantitative analyses. To test hypotheses H1-3, H2-1, and H2-3, we conducted Analysis of Variance (ANOVA). ANOVA examines the relationships between multiple nominal independent variables and a continuous dependent variable. In cases where homogeneity of variance was violated (Levene’s test resulted in a p-value below .05), we employed Welch’s ANOVA test (H1-1, H1-2, and H2-2). For hypotheses H3-1, H3-2, H3-3, H4-1, H4-2, and H4-3, independent samples t-tests were conducted. This method examines differences between two groups based on nominal independent variables and continuous-scale dependent variables.
As an additional exploration, we conducted a General Linear Model (GLM) to examine the interaction effects among our independent variables on participants’ perceived consistency, transparency, and voice. This analysis demonstrated how different combinations of component choices impact users’ perceived fairness. Additionally, we evaluated how each flagging component impacted flaggers’ cognitive burden. To test whether different choices of classification schemes, posting guidelines, and moderator type affect participants’ cognitive load, we conducted three separate ANOVA tests, one for each of those components. In the case of the text box, we conducted an independent samples t-test to see how the availability of the text box affects users’ perceived cognitive burden.
@@ 4.3.2.   Qualitative Analysis
The survey included open-ended questions to elicit suggestions for enhancing consistency, transparency, and voice in the flagging mechanism. After cleaning the data, we collected 1,741 valid responses for these three questions: Consistency (N = 657), Transparency (N = 691), and Voice (N = 393). Next, we performed an inductive analysis (Corbin and Strauss, 2015) on these responses using NVivo v.14.
Although the survey provided separate boxes for suggestions related to each aspect of fairness (consistency, transparency, and voice), respondents frequently combined and expressed their responses in a single long answer, highlighting the interconnected nature of these aspects. Thus, we integrated responses to these separate sections and analyzed them together to understand the participants’ intentions better. In addition, we excluded the responses that merely pointed out the importance of each flagging mechanism component without any additional elaboration, e.g., a response just stating ‘text box’ was excluded. This step helped us surface more nuanced insights.
The two coauthors independently analyzed the initial 20% of open-ended responses concerning consistency, transparency, and voice, and coded them (Miles et al., 2014). Some responses were so detailed and complex that we attached multiple codes to them. Subsequently, we engaged in iterative discussions to compare and refine our codes, often reflecting on emerging concepts and quoted responses, and achieving consensus on code definitions and applications through collaborative discussions. Through this process, we next combined and distilled our codes into five key themes that we present as our findings. Sec. 5.2 presents the results of our qualitative analysis.
@@ 5.1.1.   Rule Violation Classification
In this section, we examine how different implementations of rule violation classification schemes affect user perceptions. Table 5 shows the mean values of perceived consistency, transparency, and voice in participant groups that encountered the three classification schemes: absent, simple, and detailed. We also present our results on whether these perceptions significantly differ across the schemes.
Table 5. Mean Values of Perceived Consistency, Transparency, and Voice across Different Classification Scheme Conditions.
Fairness aspect Classification scheme Mean SD SE
Consistency None 5.53 1.45 .05
Simple 5.61 1.39 .05
Detailed 5.48 1.51 .05
Transparency None 5.29 1.43 .05
Simple 5.35 1.32 .04
Detailed 5.27 1.39 .04
Voice None 5.23 1.66 .05
Simple 5.37 1.45 .05
Detailed 5.30 1.52 .05
  • Total N = 2,936 (None = 968, Simple = 980, Detailed = 988)
Perceived Consistency. The mean value of perceived consistency in the simple classification scheme condition (M = 5.61) slightly exceeds that in the no (M = 5.53) or detailed (M = 5.48) classification scheme conditions. However, running Welch’s ANOVA test revealed that these differences are not significant (F(2, 1953.60) = 1.88, p = .15). Further, we compared differences between each pair of classification schemes (i.e., absent vs. simple, absent vs. detailed, and simple vs. detailed classification) by conducting t-tests. These tests revealed that participants in the simple classification scheme condition had a significantly higher perceived consistency than those in the detailed condition (t(1955.99) = 1.91, p = .01). There were no significant differences in perceived consistency between the absent and simple classification scheme conditions (t(1946) = -1.182, p = .12) or between the absent and detailed classification scheme conditions (t(1954) = .74, p = .23).
Perceived Transparency. The mean value of perceived transparency in the simple classification scheme condition (M = 5.35) slightly exceeds that in the absent (M = 5.29) and detailed (M = 5.27) conditions. However, running Welch’s ANOVA test showed that these differences are not significant (F(2, 1952.21) = 1.02, p = .36). Further, t-tests comparing differences between each pair of classification schemes revealed no significant differences in perceived transparency among any pairs.
Perceived Voice. Our descriptive analysis showed that the mean value of perceived voice in the simple classification scheme condition (M = 5.37) slightly exceeds that in the absent (M = 5.23) and detailed (M = 5.30) conditions. Again, Welch’s ANOVA test showed no significant differences among the three classification scheme conditions (F(2, 1947.25) = 1.90, p = .15). However, t-tests revealed that participants in the simple classification scheme condition had a significantly higher perceived voice than those in the absent condition (t(1905.134) = 1.94, p = .03). There were no significant differences between the absent and detailed classification scheme conditions (t(1929.78) = -.91, p = .18) or between the simple and detailed classification scheme conditions (t(1966) = 1.07, p = .14).
To summarize, we reject H1-1 given that contrary to our hypothesis, participants shown a simple classification scheme have a higher perceived consistency than those shown a detailed one. We also reject H1-2 since the different classification schemes do not significantly impact perceived transparency. However, we found partial support for H1-3, given that participants with a simple classification scheme have a higher perceived voice than those without one.
@@ 5.1.2.   Posting Guidelines
We examined how displaying posting guidelines with different granularity levels during flagging affect user perceptions. Table 6 shows the mean values of perceived consistency, transparency, and voice in participant groups that encountered posting guidelines with three distinct levels of granularity: absent, simple, and detailed. We then analyzed whether these perceptions differ significantly across the three schemes.
Table 6. Mean Values of Perceived Consistency, Transparency, and Voice across Different Guidelines Conditions.
Fairness aspect Guidelines level Mean SD SE
Consistency None 5.51 1.48 .05
Simple 5.53 1.45 .05
Detailed 5.59 1.43 .05
Transparency None 5.15 1.45 .05
Simple 5.33 1.37 .04
Detailed 5.44 1.29 .04
Voice None 5.24 1.61 .05
Simple 5.31 1.54 .05
Detailed 5.35 1.49 .05
  • Total N = 2,936 (None = 993, Simple = 971, Detailed = 972)
Perceived Consistency. The mean value of perceived consistency increased sequentially across the three conditions, being lowest in the absent guidelines condition (M = 5.51), higher in the simple guidelines condition (M = 5.53), and highest in the detailed guidelines condition (M = 5.59). However, running a Welch’s ANOVA test showed that these differences are not significant (F(2,2933) = .91, p = .40). Further, we compared differences between each pair of conditions (i.e., absent vs. simple, absent vs. detailed, and simple vs. detailed guidelines) by conducting t-tests. Test results indicated no significant differences in perceived consistency between any of the pairs.
Perceived Transparency. The mean value of perceived transparency also increased sequentially, being lowest in the absent guidelines condition (M = 5.15), higher in the simple guidelines condition (M = 5.33), and highest in the detailed guidelines condition (M = 5.44). In this case, the results of Welch’s ANOVA confirmed a significant relationship between the levels of guidelines and perceived transparency (F(2, 1952.81) = 11.36, p ¡ .001). The Games-Howell post hoc analysis showed that participants’ perceived transparency was significantly higher in the simple guidelines condition compared to the no guidelines condition (MD = 0.18, SE = 0.06, p = .01). It was also higher in the detailed guidelines condition compared to the no guidelines one (MD = 0.29, SE = 0.06, p ¡ .001). However, there was no significant difference in perceived transparency between the simple and detailed guidelines conditions (MD = -0.11, SE = 0.06, p = .15).
Perceived Voice. As in the case of perceived consistency and transparency, the mean values of perceived voice increased with increasing granularity of posting guidelines, being lowest in the absent guidelines condition (M = 5.24), higher in the simple guidelines condition (M = 5.31), and highest in the detailed guidelines condition (M = 5.35). However, an ANOVA test showed that these differences are not statistically significant (F(2, 2933) = 1.20, p = .30). Additional t-tests comparing each pair of conditions revealed no significant differences between any of the pairs.
In sum, we found that differences in granularity levels of posting guidelines do no significantly impact participant perceptions of consistency and voice, thus rejecting H2-1 and H2-3. However, we found partial support for H2-2: participants encountering simple or detailed guidelines had a higher perception of transparency than those in the absent guidelines condition.
@@ 5.1.3.   Text Box
We examined how the availability of a text box affects user perceptions. Table 7 shows the mean values of perceived consistency, transparency, and voice in participant groups that encountered the two different text box conditions: absent and present. We then conducted independent samples t-tests to evaluate how offering this text box shapes users’ perceptions.
Table 7. Mean Values of Perceived Consistency, Transparency, and Voice for the Two Text Box Conditions: Present and Absent.
Fairness aspect Text box Availability Mean SD SE
Consistency Absent 5.54 1.44 .04
Present 5.55 1.46 .04
Transparency Absent 5.27 1.38 .04
Present 5.33 1.38 .04
Voice Absent 4.94 1.64 .04
Present 5.67 1.35 .04
  • Total N = 2,936 (None = 1,492, Present = 1,444)
Perceived Consistency & Transparency. The mean value of perceived consistency in the text box condition (M = 5.55) is almost the same as in the no text box condition (M = 5.54). The mean value of perceived transparency in the text box condition (M = 5.33) slightly exceeds that in the no text box condition (M = 5.27). The t-test results indicated no statistically significant differences in perceived consistency (t(2934) = -0.077, p = .469) or transparency (t(2934) = -1.146, p = .126) between participants in the two text box conditions (absent vs. present).
Perceived Voice. The mean value of perceived voice in the text box condition (M = 5.67) surpasses that in the no text box condition (M = 4.94). A t-test confirmed that this difference across the two conditions is significant (t(2859.50) = -13.32, p ¡.001). Specifically, users provided with a text box (M = 5.67, SD = 1.35, SE = .04) reported significantly higher perceived voice compared to those without one (M = 4.94, SD = 1.64, SE = .04).
To sum up, our analysis shows that the availability of open-ended text boxes that let users describe their reasons for flagging does not significantly affect their perceptions of consistency and transparency, leading us to reject H3-1 and H3-2. However, we found support for H3-3: participants in the text box condition reported a higher perceived voice compared to those without such a box.

### v2

@@ 3.4. Moderator Type
4. Methods
We begin this section by describing how we implemented 54 distinct flagging scenarios using Qualtrics to survey our participants. Next, we offer information about our survey deployment and detail our ethical considerations. Finally, we describe the methods we used to analyze the Likert scale and open-ended responses from our survey data. We preregistered this study at OSF. [^4] https://osf.io/982wd/?view_only=91e6792f777b44b582fd6bb480ab142c
4.1. Survey Design
Our mixed methods study deployed a triangulation design (Ivankova and Creswell, 2009) in which we collected both quantitative and qualitative data using a questionnaire that contained close-ended and open-ended questions. Both methods were given equal weight (Ivankova and Creswell, 2009) and the mixing of the two methods occurred during the interpretation of the study results (Section 6). This approach allowed us to compare and contrast our findings to produce well-validated conclusions and gain an in-depth understanding of how the overall design and procedural space of flag mechanisms shape flaggers’ fairness attitudes.
To collect data for this study, we created an online survey using Qualtrics software and distributed it via Lucid Theorem, an academically-oriented survey platform, to collect responses from U.S. adult internet users. The survey began with a page displaying seven examples of rule-violating posts, shown in Table 4. Warning: The example posts in Table 4 may be disturbing to some readers.
No. Post
     6 Omg just got tons of Bucks from here! - [xyz phishing site].com.
The flag submission choices shown to participants followed a between-subjects factorial design created by combining different implementation choices for the four flagging components. It comprised 3 (Rule violation classification schemes) $\times$ 3 (Granularity of posting guidelines) $\times$ 2 (Availability of a text box) $\times$ 3 (Types of moderators) different conditions. Depending on their assigned scenario, participants experienced variations in the design of the rule violation classification scheme, the level of detail regarding platform guidelines, the presence or absence of an open-ended text box, and information about the moderator (human, bot, or left unspecified).
@@ 4.1.1.   Flagging Scenarios
Section 3 details how we operationalized the four components of flagging designs within the Qualtrics software. In summary, Figure 4 illustrates the flagging components and an example of flagging design flow a participant could have encountered in the survey. The left column enumerates each component and its different levels (i.e., conditions). The right column shows one example of the flagging scenario a participant might have encountered. Once a participant selected one of the seven examples of inappropriate posts listed above, the flagging process (scenario) started. In this particular scenario, a participant would have been required to choose one of the rule violations from the primary classification scheme for the post they selected. Then, the participant would have been offered a text box to describe the rule violation in detail. The final step would have been a message of gratitude for flagging, informing them that an auto-moderator will review the post they flagged.
@@ 4.1.2.   Dependent Variables: Procedural Fairness
We used responses to these questions to operationalize participants’ perceptions of consistency, transparency, and voice, respectively. Participants indicated their responses to these questions on a 7-point Likert scale that ranged from Strongly Disagree to Strongly Agree.
Additionally, to answer RQ 4, the survey featured three open-ended questions asking participants to provide suggestions on how the flagging process could be further enhanced. These questions were:
@@ 4.1.3.   Usability
    “How demanding was it for you to use this flagging mechanism to report the post you selected?” (rated on the scale of ‘Very demanding’ to ‘Very undemanding’)
@@
    “How likely are you to use this flagging mechanism to report any inappropriate post on a social media platform you use?” (rated on the scale of ‘Very likely’ to ‘Very unlikely’.)
@@ 4.2. Deployment of the Survey
Our study was considered exempt from review by our institute’s [^5] We will reveal the institute name after the peer review is completed. IRB on November 16th, 2023. Before launching the survey, we conducted a pilot survey from May 1 to May 8 to gather feedback on the survey’s wording and organization. Based on this feedback, we revised the survey questionnaire by adjusting certain words and adding signposts. After this, we rolled out our main survey, which targeted U.S. adult internet users aged 18 and above. Participants were recruited through Lucid Theorem (https://lucidtheorem.com), a survey company providing access to nationally representative samples, from May 16 to May 18, 2024. Participants who completed the survey received \$1.50 in compensation via the Lucid system. In total, we collected responses from 3,650 participants.
We excluded respondents who opted out during the survey (N = 496), [^6] Our review of collected data shows that 253 participants opted out on the first page where we requested participation consent, 7 dropped out at the next question requesting selection of a post that participants would like to flag, 148 dropped out in the flag submission stage, and 88 dropped out after answering the flag submission questions. We found no noticeable trends in the number of dropouts across the different flag implementation scenarios. spent less than 1 minute or more than 50 minutes on the survey (N = 40), and exhibited straight-lining behavior by selecting the same response for all questions (N = 178). Following these pre-processing steps, we retained 2,936 survey responses, which we used for our subsequent analyses. Appendix A describes the demographic distribution of our survey sample. On average, the participants took 393.2 seconds to complete the survey. As mentioned above, the survey began with participants selecting a post to flag from the list of rule-violating posts presented in Table 4. Table 9 (Appendix B) shows the distribution of post choices made by survey participants and how average fairness perceptions vary across these choices — we did not find any noticeable trends in fairness perceptions based on participants’ post choices.
@@ 4.3.1.   Quantitative Analysis
We used SPSS Version 29 to analyze our quantitative data. Sec. 5.1 presents the results of our quantitative analyses. To test hypotheses H1-3, H2-1, and H2-3, we conducted one-way Analysis of Variance (ANOVA) tests. ANOVA examines whether statistically significant differences exist between three or more groups. In cases where homogeneity of variance was violated (Levene’s test resulted in a p-value below .05), we employed Welch’s ANOVA test (H1-1, H1-2, and H2-2). For testing hypotheses H3-1, H3-2, H3-3, H4-1, H4-2, and H4-3, we conducted independent samples t-tests. This method examines whether statistically significant differences exist between two groups.
As an additional exploration, we built a General Linear Model (GLM) to examine the interaction effects among our independent variables on participants’ perceived consistency, transparency, and voice. This analysis demonstrated how different combinations of component choices impact users’ perceived fairness. Additionally, we evaluated how each flagging component impacted flaggers’ cognitive burden. To test whether different choices of classification schemes, posting guidelines, and moderator type affect participants’ cognitive load, we conducted three separate ANOVA tests, one for each of those components. In the case of the text box, we conducted an independent samples t-test to see how the availability of the text box affects users’ perceived cognitive burden.
@@ 4.3.2.   Qualitative Analysis
The survey included open-ended questions (Sec. 4.1.2) to elicit suggestions for enhancing consistency, transparency, and voice in the flagging mechanism. After cleaning the data, we collected 1,741 valid responses for these three questions: Consistency (N = 657), Transparency (N = 691), and Voice (N = 393). Next, we performed an inductive analysis (Corbin and Strauss, 2015) on these responses using NVivo v.14.
Although the survey provided separate boxes for suggestions related to each aspect of fairness (consistency, transparency, and voice), respondents frequently combined and expressed their responses in a single long answer, highlighting the interconnected nature of these aspects. Thus, we integrated responses to these separate sections and analyzed them together to understand the participants’ intentions better. In addition, we excluded the responses that merely pointed out the importance of each flagging component without any additional elaboration, e.g., a response just stating ‘text box’ was excluded. This step helped us surface more nuanced insights.
Next, the two coauthors independently analyzed the initial 20% of open-ended responses concerning consistency, transparency, and voice, and coded them. Some responses were so detailed and complex that we attached multiple codes to them. Subsequently, we engaged in iterative discussions to compare and refine our codes, often reflecting on emerging concepts and quoted responses, and achieving consensus on code definitions and applications through collaborative discussions. After coding the entire dataset, we refocused our analysis at the broader level of themes rather than codes. Specifically, we iteratively sorted related codes into potential themes by using mind-maps and attending to the relationships between codes and between emerging themes (Braun and Clarke, 2006). Through this process, we next reviewed and refined our candidate themes and arrived at five key themes we present as our findings. Sec. 5.2 presents the results of our qualitative analysis.
@@ 5.1. Quantitative Findings
Table 5. Mean Values of Perceived Consistency, Transparency, and Voice Across Different Conditions of Flagging Components
(Likert Scale: Strongly Disagree = 1 to Strongly Agree = 7). Standard deviation (SD) values are in parentheses. N represents the number of participants placed in the condition.
     Condition Group Condition N Consistency (M, SD) Transparency (M, SD) Voice (M, SD)
     Classification Scheme None 968 5.53 (1.45) 5.29 (1.43) 5.23 (1.66)
Simple 980 5.61 (1.39) 5.35 (1.32) 5.37 (1.45)
Detailed 988 5.48 (1.51) 5.27 (1.39) 5.30 (1.52)
     Guidelines Level None 993 5.51 (1.48) 5.15 (1.45) 5.24 (1.61)
Simple 971 5.53 (1.45) 5.33 (1.37) 5.31 (1.54)
Detailed 972 5.59 (1.43) 5.44 (1.29) 5.35 (1.49)
     Text Box Availability Absent 1,492 5.54 (1.44) 5.27 (1.38) 4.94 (1.64)
Present 1,444 5.55 (1.46) 5.33 (1.38) 5.67 (1.35)
     Moderator Type Not Available 988 5.56(1.46) 5.32 (1.38) 5.33 (1.54)
Human 972 5.52(1.45) 5.27(1.41) 5.26 (1.55)
Bot 976 5.55 (1.45) 5.33(1.34) 5.32(1.55)
Table 5 shows the mean values of perceived consistency, transparency, and voice reported by participant groups that encountered different conditions for the four flagging components. Given that we ran three tests for each flagging component, we estimated the statistical significance of each result following Bonferroni correction ($\alpha$05/3). Table 6 summarizes the results for our hypothesis.
Table 6. Summary of Research Hypotheses, Variables, and Outcomes
Hypothesis Independent Variable Dependent Variable Outcome
H1-1 More granular classification schemes will increase perceived consistency. Granularity of classification scheme Consistency Not supported
H1-2 More granular classification schemes will increase perceived transparency. Transparency Not supported
H1-3 More granular classification schemes will increase perceived voice. Voice Not supported
H2-1 More granular posting guidelines will increase perceived consistency. Granularity of posting guidelines Consistency Not supported
H2-2 More granular posting guidelines will increase perceived transparency. Transparency Partially supported
H2-3 More granular posting guidelines will decrease perceived voice. Voice Not supported
H3-1 Offering a text box will decrease perceived consistency. Text box availability Consistency Not supported
H3-2 Offering a text box will increase perceived transparency. Transparency Not supported
H3-3 Offering a text box will increase perceived voice. Voice Supported
H4-1 A bot moderator will have higher perceived consistency than an unknown or human moderator. Moderator type Consistency Not supported
H4-2 A bot or human moderator will have higher perceived transparency than an unknown moderator. Transparency Not supported
H4-3 A human moderator will have higher perceived voice than a bot or unknown moderator. Voice Not supported
@@ 5.1.1.   Rule Violation Classification
We tested Hypotheses 1 by examining how different implementations of rule violation classification schemes affect user perceptions through ANOVA. First, we rejected H1-1 given that contrary to our hypothesis, participants shown a simple classification scheme have a higher perceived consistency than those shown a detailed one. Further, no significant differences were observed between the absent and simple conditions, nor between the absent and detailed conditions. We also rejected H1-2 since the different classification schemes do not significantly impact perceived transparency. Finally, we rejected H1-3, given that differences in classification scheme do not significantly impact perceived voice.
@@ 5.1.2.   Posting Guidelines
To test Hypotheses 2, we used ANOVA tests to examine how displaying posting guidelines with different granularity levels affect user perceptions. We found that these differences do not significantly impact participant perceptions of consistency and voice, thus rejecting H2-1 and H2-3. However, we found partial support for H2-2 — participants encountering simple (M = 5.33) or detailed guidelines (M = 5.44) have a significantly higher perception of transparency than those in the absent guidelines condition (M = 5.15, F(2, 1952.81) = 11.36, p ¡ .001).
@@ 5.1.3.   Text Box
We conducted t-tests to examine how the availability of a text box affects user perceptions to test Hypotheses 3. Our analysis shows that the availability of open-ended text boxes that let users describe their reasons for flagging does not significantly affect their perceptions of consistency and transparency, leading us to reject H3-1 and H3-2. However, we found support for H3-3 — participants in the text box condition (M = 5.67) report a significantly higher perceived voice compared to those without such a box (M = 4.94, t(2859.50) = -13.32, p ¡.001).